if P(A)=1/3 and P(B)=1/6 . What is P(A U B) and P (A ∩ B) ?
Answer #1

Assuming A and B are independent,

P(An B) = P(A) * P(B) =-+-- 3 6 18

and

P(AU B) = P(A) + P(B) _ P(An B)

3 6 18

6+3-1 P(AUB)

P(AUB) 18

Rightarrow P(Acup B) = rac{4}{9}

But,

If A and B are not independent, then it is not possible to compute P(Acap B) , and hence P(A U B)